5) Investment E pays $250 at the beginning of every year for the next 10 years (a total of 10 payments). Question 25 (3 points) A couple decides to purchase a beach condominium in Hilton Head Island, South Carolina. The agreed purchase price will be $298,000. The couple will pay 20% down on the condo and finance the remaining balance with a 30-year mortgage. The terms of the mortgage are 4.80% APR (with monthly compounding). The couple will inherit a trust fund exactly ten years from the purchase of the house. The fund is expected to be redeemed for $219,500. If the couple will pay off the remaining balance on the condo at that point, what will be the money left in the trust after the payment? Format $12.344 as 12344 Your Answer: Answer Question 26 13 points) A grandfather tells his grandson that he has created a trust fund for him. The terms of the trust are as follows. $9.800 per year for 30 years. The money will be invested in an account does the grandfather need to endow the trust with today to fund ail the payments? Format that pays 4% annual interest. The payments wit start 15 years from today. How much money $12.345 as 12345 Your Answer f7 f8 f9 f10 f11 15 f6
Answer #1
QUESTION 25:
Amount of finance required = 298000*(1-20%) = $          2,38,400
Monthly installments (using the loan amortization formula) = 238400*0.004*1.004^360/(1.004^360-1) = $          1,250.80
Balance outstanding at the end of 10 years = PV of the remaining 240 installments = 1250.80*(1.004^240-1)/(0.004*1.004^240) = $    1,92,739.90
Amount left from the money received from the trust after paying off the balance of the loan = 219500-192739.90 = $       26,760.10
QUESTION 26:
The amount to be accumulated in the fund at EOY 15 = 9800*(1.04^30-1)/(0.04*1.04^30) = $    1,69,461.93
Amount to be endowed today = 169461.93/1.04^15 = $       94,096.19